OpenStudy (anonymous):

I have another order pair.. anyone mind helping me out? 4x+y=0 8x-y=6

jimthompson5910 (jim_thompson5910):

you can add the equations to get 4x+y=0 8x-y=6 -------- 12x + 0y = 6 so 12x = 6

jimthompson5910 (jim_thompson5910):

if 12x = 6, then what is x ?

jimthompson5910 (jim_thompson5910):

you'll notice that with elimination and substitution, the goal is to get rid of one variable so you can solve for the other

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

no, but close

jimthompson5910 (jim_thompson5910):

you would divide both sides by 12 to isolate x

OpenStudy (anonymous):

Oh, okay so I'll have to divide 12 from 6 right?

jimthompson5910 (jim_thompson5910):

yes 12x = 6 x = 6/12 x = 1/2

jimthompson5910 (jim_thompson5910):

now use that to find y

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

:O

jimthompson5910 (jim_thompson5910):

pick any equation you want and plug x = 1/2 into it

jimthompson5910 (jim_thompson5910):

so let's pick the first equation

jimthompson5910 (jim_thompson5910):

4x+y=0 4*(1/2)+y=0 4/2 + y = 0 2 + y = 0 y = ??

OpenStudy (anonymous):

Hmm, I'm a bit confused how would I answer this? :/

jimthompson5910 (jim_thompson5910):

if 2+y = 0, then what is y?

OpenStudy (anonymous):

would it be a negative number?

jimthompson5910 (jim_thompson5910):

yes it would be

OpenStudy (anonymous):

ahhh, sorry i'm horrible at math!

jimthompson5910 (jim_thompson5910):

you're ok, what do you get for y?

jimthompson5910 (jim_thompson5910):

and this is why you practice practice practice

OpenStudy (anonymous):

if you want to get rid of a number from one side add the opposite of it to both sides of the equation...

jimthompson5910 (jim_thompson5910):

if it helps at all, think of 2+y as y+2

OpenStudy (anonymous):

sorry, i'm confused..... :(

jimthompson5910 (jim_thompson5910):

2+y = 0 y+2 = 0 how do you undo that +2

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

yep you subtract 2 from both sides

jimthompson5910 (jim_thompson5910):

to get y = -2

OpenStudy (anonymous):

so, the answer for y will be -2?

jimthompson5910 (jim_thompson5910):

correct, the ordered pair is (1/2, -2)

OpenStudy (anonymous):

Thank you so much! but one more question can I reduce the fraction or is that the lowest it'll be lol just curious

jimthompson5910 (jim_thompson5910):

1/2 is as reduced as it gets

OpenStudy (anonymous):

Oh, okay! Thank you so much for the help!!!

jimthompson5910 (jim_thompson5910):

sure thing